OpenStudy (anonymous):

Find the integral: (x+2)(x-3)^1/2

jhonyy9 (jhonyy9):

what mean the exponent 1/2 ?

OpenStudy (anonymous):

sqrt

jhonyy9 (jhonyy9):

yes right so than how will be your exercise ?

OpenStudy (anonymous):

Dint getcha!!

OpenStudy (callisto):

\[\int(x-2)(x-3)^{\frac{1}{2}}dx\] Let u = x-3 du = ... ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

dx

OpenStudy (callisto):

Hmm... du = (1) dx So, du = dx x-2 = (x-3)+1 = u+1 So, the integral becomes \[\int (u+1)u^{\frac{1}{2}} du\] Can you find the integral now?

OpenStudy (anonymous):

X-2 is raised to power sqrt(1/2)

OpenStudy (callisto):

Is the question (a)\[\int(x-2)(x-3)^{\frac{1}{2}}dx\] Or (b) \[\int[(x-2)(x-3)]^{\frac{1}{2}}dx\]?

OpenStudy (anonymous):

Integral of { (x+2)*((x-3)^1/2)}

OpenStudy (anonymous):

@cambrige Expand \(\bf u^{1/2}(u+1)\) and then integrate.

OpenStudy (anonymous):

how to expand u^1/2???

OpenStudy (callisto):

x+2 = x-3 + 3 +2 = (x-3) + 5 = u+5 So, the integral becomes \[\int (u+5)u^\frac{1}{2}du\]

OpenStudy (anonymous):

@Callisto Where did the u + 5 come from. Isn't it supposed to be u + 1?

OpenStudy (callisto):

The asker typed the question wrong. Check his/her previous comment.

OpenStudy (anonymous):

still not gtin

OpenStudy (callisto):

Which part?

OpenStudy (anonymous):

doent match with ans

OpenStudy (anonymous):

*doesnt

OpenStudy (callisto):

Have you integrated it?

OpenStudy (anonymous):

ans_ {2(3x+16)((x-3)^3/2)}/16+c

OpenStudy (anonymous):

yes

OpenStudy (callisto):

Would you mind showing your work?

OpenStudy (anonymous):

wait a sec

OpenStudy (anonymous):

1 min\

OpenStudy (anonymous):

I dont even know if its correct!!

OpenStudy (callisto):

Expand \((u+5)*u^\frac{1}{2}\), can you?

OpenStudy (anonymous):

tryin now

OpenStudy (callisto):

By distributive property

OpenStudy (callisto):

Calm... What do you get when you expand (a+2)a?

OpenStudy (anonymous):

a^2+2a

OpenStudy (callisto):

Yes. Same thing here. You need to expand \((u+5)u^\frac{1}{2}\) first. What do you get when you expand \((u+5)u^\frac{1}{2}\)?

OpenStudy (anonymous):

{u^2+5u}/u^1/2

OpenStudy (callisto):

That is \(u^\frac{3}{2} + 5u^\frac{1}{2}\) Now, can you integrate \(u^\frac{3}{2} + 5u^\frac{1}{2}\) with respect to u, that is can you solve \[\int u^\frac{3}{2} + 5u^\frac{1}{2}du\]?

OpenStudy (callisto):

I don't know how you got the first step.

OpenStudy (callisto):

No. You don't have to change the variable now. You change the variable after you finish doing \(\int u^\frac{3}{2} + 5u^\frac{1}{2}du\) What have you got when you do \(\int u^\frac{3}{2} + 5u^\frac{1}{2}du\)?

OpenStudy (callisto):

in terms of u.

OpenStudy (anonymous):

why to integrate again??

OpenStudy (callisto):

Check the second term. Something wrong with that.

OpenStudy (callisto):

Also, after you integrate it, remove the integral sign and "du".

OpenStudy (anonymous):

ya....sry bout dat!!.....nothing wrong with 2nd term!

OpenStudy (callisto):

Sorry, my mistake about the second term. So, we get \[\frac{2u^\frac{5}{2}}{5} + \frac{10u^\frac{3}{2}}{3}+C\] Next, replace all u by x-3. Then we're done.

OpenStudy (anonymous):

tryin

OpenStudy (anonymous):

matches with the ans